OpenStudy (anonymous):

A car accelerates for 5 seconds at 2.1ms, if the car initially had a speed of 10 ms what is the new speed

OpenStudy (anonymous):

v2 = v1 + at, where v2 is the new speed, v1 is the initial speed, a is the acceleration and t is time. All standard SI units. Thus: v2 = 10 + 2.1(5) = 20.5